nolegirl24:

Choose the correct product of (5x − 11)^2. 25x^2 − 110x + 121 25x^2 − 121 25x^2 + 121 25x^2 + 110x + 121

nolegirl24:

@dude

dude:

re-write it as \( (5x − 11) (5x − 11)\) it might help looking at it easier

dude:

Now distribute

nolegirl24:

25x^2 + 121?

nolegirl24:

That would be the answer right?

dude:

Not quite, remember -11 times 5x is -55x -55x-55x is -110x (They are both negatives)

nolegirl24:

I thought you couldnt do 5x * -11

dude:

When adding yes, but you are distributing here

dude:

Multiplying doesnt follow that

nolegirl24:

So i would have to do 5x * 5x, 5x * -11, -11 * 5x, and -11 * -11?

dude:

Yes

nolegirl24:

okay one sec

nolegirl24:

So it would be 5x * 5x = 25x^2 5x * -11 = -55x -11 * 5x = -55x -11 * -11 = 121

dude:

Yes

nolegirl24:

And then it would turn out to be like 25x^2 - 55x - 55x + 121?

dude:

Right, so now add -55x and -55x

nolegirl24:

-110 Sp it would be 25x^2 - 110x + 121?

dude:

Correct

nolegirl24:

So answer is A ^~^

dude:

Yesh :P

nolegirl24:

YAY